MULTIVARIATE PERCENTILE TESTS FOR INCOMPLETE DATA

HYO-IL PARK

Department of Statistics, Chong-ju University, Chong-ju, Choong-book, 360-764, Korea, e-mail: hipark@chongju.ac.kr

(Received May 23, 2001; revised October 19, 2001)

Abstract. In this paper, we consider the percentile test procedures for multivariate and right censored data. Because of the involvement of censoring distribution into the distribution of the proposed test statistic, we study the asymptotic normality using the estimated covariance matrix. Finally, we derive the asymptotic relative efficiency and illustrate our procedures with an example.

Key words and phrases: Asymptotic relative efficiency, noncentrality parameter, Pitman translation parameter, two sample problem.